# Quantum computing with atomic qubit arrays: confronting the cost of connectivity

### v1

@@ III Rydberg gate fidelity
The original proposals for entangling pairs of atoms and preparing many-body entanglement using the Rydberg blockade mechanism[79, 17] were introduced 25 years ago. The basic ideas and approaches that have been developed for encoding qubits in neutral atoms, their interactions, and mechanisms for performing and characterizing Rydberg quantum gates have been reviewed in many articles [80, 81, 82, 18, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96]. First observations of the blockade phenomenon in a many atom ensemble[97] were followed by demonstration of blockade between pairs of individual atoms[98, 99]. Shortly thereafter the first entanglement and two-qubit quantum gates were achieved using Rydberg interactions[100, 101, 102]. The first gate demonstrations had relatively low fidelity and it took some years for the performance to be improved as can be seen in Fig. 4. The improvement can be ascribed to new developments in several directions which together have resulted in recent rapid progress reaching fidelity well over $99\%$ in several experiments. Notably these high fidelity results were achieved with four different atomic elements 87Rb[103], 133Cs[39], 88Sr[104], and 171Yb[105, 106]. While a high fidelity two-qubit gate forms the basis for scaling to larger circuits, in the context of optimizing performance for quantum error correction the two-qubit gate fidelity may not directly translate to the fidelity that is achieved for a syndrome extraction protocol or logical operation. In other words optimizing the gate protocol using the logical performance as a cost function may lead to better performance than optimizing the two-qubit gate in isolation. Several recent papers have addressed this question and predicted improved performance for surface code syndrome extraction[107, 108, 109].
@@
Figure 4: Experimental results for Rydberg $\sf CZ$ gate fidelity. The atomic species are Rb (red circles), Cs (blue squares), Sr (purple triangles), Yb (green diamonds). References: a[100], b[101], c[102], d[110], e[111], f[112], g[113], h[114], i[115], j[116], k[117], l[44], m[118], n[119], o[103], p[70], q[120], r[39], s[49], t[48], u[106], v[121], w[105], x[40], y[104].
@@
One area of progress has been development of improved gate protocols that reduce the amount of time spent in the Rydberg state, thereby decreasing scattering errors, and which are more robust against experimental imperfections[122, 107, 123, 124]. These time-optimal gates can be designed to work well also without strong blockade[125], in which case there is some population of doubly-excited Rydberg states during the gate. In this case additional errors arise from inter-atomic forces which must be accounted for[126].
@@
Another factor contributing to improved performance has been the preparation of colder atoms. Early experiments[100, 101, 102] had atom temperatures greater than $50~{}\mu\rm K$ whereas the recent, high fidelity results were all below $10~{}\mu\rm K$. Although the Rydberg blockade interaction is intrinsically insensitive to atom temperature there are higher order effects which lead to errors that grow with temperature. In order to prepare entanglement the atoms must be in a superposition of ground and Rydberg states for some amount of time during the gate sequence. Atomic motion, due either to finite temperature, or photon recoil from the laser excitation[126], leads to a stochastic gate phase as was first pointed out in [100], which degrades the entanglement fidelity. Furthermore as the temperature increases the trapped atoms become progressively more delocalized which leads to fluctuations in the Rydberg interaction strength, as well as excitation errors due to uncertain overlap with focused Rydberg beams. This latter effect can be reduced, but not eliminated, using shaped beams with a flattened intensity profile[127, 128, 103, 40].
@@
A technical aspect of the improvement in gate fidelity that deserves highlighting is the role played by laser noise. Even when the atoms are in their motional ground state phase and intensity noise on the excitation lasers contribute to imperfect Rabi oscillations between ground and Rydberg states. Although the issue of laser noise was known in the trapped ion and clock communities[129, 130, 131] the specific impact on Rydberg gates with Rabi frequencies in the MHz range had not been studied in detail until the work of Ref. [132]. There it was shown by numerical simulations that phase noise at the frequency of the ground-Rydberg Rabi frequency was particularly detrimental. Additional analysis established limits on the spectral density of the noise for reaching a desired gate fidelity[133, 134]. These limits are not met by the external cavity diode lasers that were typically used for early Rydberg excitation experiments. It was shown convincingly in [113] that cavity filtering of the laser phase noise spectrum led to improved control fidelity and recent demonstration of “spin-locking" diagnostics has enabled direct characaterization of the errors induced by laser noise[104].
@@
With all of these advances the state of the art in Rydberg mediated $\sf CZ$ gate fidelity has reached ${\mathcal{F}}=0.9971(5)$[104]. It is likely that with further experimental improvements ${\mathcal{F}}>0.999$ will soon be achieved. This may occur first in the alkaline earth atoms Sr and Yb which use a one-photon Rydberg excitation pulse starting from the metastable 3P0 clock state. For the heavy alkali atoms Rb and Cs two-photon excitation from the ground state has been the most widely used approach, although there is also progress on one-photon Rydberg excitation of Cs for implementing an adiabatic dressing gate[135, 121]. With two-photon excitation, high fidelity requires extremely low intensity noise in order to stabilize the differential AC Stark shift between ground and Rydberg states[110] that is significant when operating at high intensity and large detuning from the intermediate state used for ladder excitation. High intensity and large detuning are requirements in order to have a fast Rabi rate to minimize Rydberg scattering errors, while also minimizing scattering from the intermediate state.
@@
[…]
It was shown in [136] assuming the atomic qubits start in a separable pure state, and considering the growth of the von Neumann entropy during the gate, that in order to create one unit of entanglement we require
@@
It is instructive to estimate this bound using atomic parameters and compare with known gate protocols. As an example consider two Cs atoms excited to the $|68p_{3/2},m_{j}=3/2\rangle$ state that are separated by $R=2.5~{}\mu\rm m$ with the interatomic axis aligned along the quantization axis. The Rydberg intgeraction energy computed with the Pair interaction code[137] is $V=2\pi\times 215~{}\rm MHz$ and the Rydberg lifetime in a room temperature environment is[138, 139] $\tau_{\rm R}=198~{}\mu\rm s$. These values give
@@
There is an important open question of how to design $\Omega(t)$ to reach this bound. The original Rydberg $\pi-2\pi-\pi$ gate proposal[79] has an error scaling[140] $\epsilon\sim 1/(V\tau_{\rm R})^{2/3}$ which is found from optimizing the Rabi frequency to balance errors between imperfect rotations at finite interaction strength ( $\sim|\Omega/V|^{2}$) and the Rydberg scattering error $(\sim 1/\Omega\tau_{\rm R}).$ It turns out that the $\pi-2\pi-\pi$ gate can be modified to have an error scaling as $1/(V\tau_{\rm R})$[141], and also other protocols that rely on individual addressing of control and target qubits reach the optimum $1/(V\tau_{\rm R})$ scaling[142].
@@
Nevertheless, protocols that have $\Omega_{1}(t)\neq\Omega_{2}(t)$ so the temporal waveforms are different on the two qubits require more complex control systems. Adiabatic proposals for gates with symmetric qubit control $(\Omega_{1}(t)=\Omega_{2}(t))$ have been proposed[143, 144], and the qubit symmetric, two-pulse protocol from[115] was the first to demonstrate fidelity above 97%. The time-optimal protocols[122, 123, 124] which have delivered the highest gate fidelities are qubit symmetric but do not rely on adiabaticity to return all initial states to the ground state at the end of the gate. Instead they rely on fast driving to minimize time spent in the Rydberg state with a time dependent phase profile that ensures that initial states with both one- and two-atoms Rydberg coupled completely return to the ground state.
@@
In the limit of strong blockade the protocols that come closest to the limit set by Eq. (5) are the two-atom dark state gate[142] which reaches $\epsilon/\epsilon_{\rm min}\simeq 19$, the time-optimal gate[122] which reaches $\epsilon/\epsilon_{\rm min}\simeq 15$ and the modified $\pi-2\pi-\pi$ gate[141]. With strong blockade additional errors due to simultaneous Rydberg excitation of both atoms, which leads to inter-atomic forces, are negligible. In the limit of weak blockade modified time-optimal profiles were found to reach $\epsilon/\epsilon_{\rm min}=2.1$[125]. Including two-atom interaction dephasing and the effect of photon recoils the predicted error increases to $\epsilon/\epsilon_{\rm min}=3.0$. Photon recoil heating can be minimized using more complex time-dependent pulses[145, 146, 147], but this has not yet been demonstrated in practice with Rydberg gates.
IV The connectivity challenge
@@
Running complex quantum algorithms on a digital quantum computer requires gate operations involving qubits that may not be next to each other. Any physical interaction between qubits has a finite range. While there are settings for which all qubits are interconnected, such cases do not support scaling to arbitrarily large qubit numbers. As an example consider coupling many atomic qubits to the mode of an optical resonator such that photons mediate all to all connectivity[148, 149, 150]. Even though photon mediated coupling does not intrinsically have a limited range there is a practical limit to how many atoms can be strongly coupled to a single cavity.
@@
For quantum computation without error correction the additional gate operations needed to connect non-proximal qubits add operational overhead and reduce the useful computational depth. The severity of establishing connectivity depends on the algorithm being run, and can be reduced with optimized compilation that is aware of the physical level hardware capabilities[151, 152]. There are also encoding techniques for specific problems that effectively provide long-range interactions at the cost of consuming more physical qubits[153, 154]. A well designed error corrected processor that operates with logical qubits should have a computational depth that greatly exceeds that of a NISQ machine. Nevertheless, useful error correction does not imply an arbitrarily small logical error rate, just an error rate that is smaller than that of the underlying physical operations. Thus also logical processors will suffer reduced useful circuit depth due to resources required for connectivity, as well as longer operation times.
@@
$p_{\rm L}=0.08(p/p_{\rm th})^{0.58d-0.27}.$ (7)
where the threshold physical error rate is $p_{\rm th}=0.0053.$ Taking ${\mathcal{F}}=0.9992$ or $p=0.0008$, which appears within reach of Rydberg gates, a code distance of $d=10$ is sufficient to reach within about a factor of two of $1/p_{\rm L}=10^{6}$. At this code distance 100 logical qubits would require 19900 physical qubits, plus the overhead of magic state distillation for non Clifford operations. Such a “Megaquop" machine with 100 error corrected logical qubits is likely to be of great interest from a scientific perspective[157]. While atom arrays are currently one of the most promising approaches for reaching this combination of low error rates and high qubit counts the clock rate of neutral atoms is slow compared to other platforms. With that challenge in mind the following sections provide a comparison of logical gate times based on several different approaches.
IV.1 Long-range Rydberg interactions
The rotated surface code has $[[n,k,d]]=[[d^{2},1,d]]$ and the gate time for a transversal logical $\sf CZ$ using sequential long-range Rydberg interactions is $t_{\rm gate}\leq d^{2}(t_{\sf CZ}+t_{\rm beam})$ where $t_{\sf CZ}$ is the time for the Rydberg gate and $t_{\rm beam}$ is the time to point the optical control beams at a new atom pair. This is a strict equality for sequential gate operations and can be quicker if some of the atom pairs are far enough separated to allow for parallel gate operations (see [125] for an analysis of the separation required to minimize interference effects between gate pairs).
Figure 6 shows the logical gate time assuming $d=10$. The calculated times which can be well under $100~{}\mu\rm s$ assume two-qubit Rydberg gates with an optical control system that excites pairs of atoms using dynamically scanned, tightly focused beams[44, 39]. The achievable gate speed is strongly dependent on the performance of the optical beam scanning device. Acousto-optic deflectors are compatible with the sub-microsecond times assumed in the figure, whereas other technologies such as spatial light modulators and digital mirror arrays operate at much slower speeds. The combination of these approaches may allow fast reconfiguration together with the ability to address large qubit arrays[158].
A limitation of this approach is the finite range of the Rydberg interaction. At $d=10$ and a tight spacing between physical qubits of $r=1.4~{}\mu\rm m$ the required interaction distance is $R=dr=14~{}\mu\rm m$ which is beyond the range of demonstrated Rydberg gates. A high fidelity gate at this distance appears feasible by combining progress in several directions. The use of highly excited Rydberg states can provide close to $V=2\pi\times 6~{}\rm MHz$ of interaction strength at this distance as shown in Fig. 7. With this value for $V$ a modified time-optimal gate profile that provides high fidelity with modest interaction strength can reach ${\mathcal{F}}=0.9992$ with a gate time of $0.46~{}\mu\rm s$[125]. With an optical beam scanning time of $0.5~{}\mu\rm s$ this leads to a logical gate time of about $90~{}\mu\rm s$.
Extending this approach to larger code distance which implies larger $R$ appears challenging. Our assumption of $r=1.4~{}\mu\rm m$ atomic spacing cannot be significantly reduced without running in to the limit set by the wavefunction of a Rydberg excited electron overlapping with a neighboring ground state atom. The Rydberg interaction strength scales as $1/R^{3}$ at short distances (limit of resonant dipolar interactions) and $1/R^{6}$ at long-range(van der Waals limit) [159]. Judicious application of dc or microwave fields can be used to reduce the Förster defects and extend the $1/R^{3}$ scaling to longer distances[160, 161, 162, 163].
Figure 7: Rydberg interaction strength for Cs atoms excited to $2100s_{1/2},m_{j}=1/$ 2 states quantized along the atomic separation axis ($\hat{z}$) with a magnetic field of $B_{z}=10^{-5}~{}\rm T$ Calculation performed with pair interaction code[137].
It is perhaps interesting to check how large the interaction could ideally be at say $R=20~{}\mu\rm m$ which could enable $d=15$ and 100 Megaquop performance. At Förster resonance the interaction of two atoms excited to Rydberg states that interact through a single channel can be expressed as $V(R)=-C_{3}D^{1/2}/R^{3}$ where $D$ is an angular factor depending on the interaction channel and
$C_{3}=\frac{e^{2}}{4\pi\epsilon_{0}}\frac{\langle\gamma_{\alpha}||r_{a}||\gamma_{a}\rangle\langle\gamma_{\beta}||r_{b}||\gamma_{b}\rangle}{\sqrt{2j_{\alpha}+1}\sqrt{2j_{\beta}+1}},$ (8)
@@ IV.1 Long-range Rydberg interactions
where $e$ is the electronic charge, $\epsilon_{0}$ is the permittivity of free space, the reduced matrix elements are calculated in the fine structure basis, and $\gamma$ denotes the quantum numbers of the initial Rydberg levels $a,b$ that are coupled to levels $\alpha,\beta$ through a Förster interaction[159, 164]. Consider as an example Cs $a,b$ states $n_{a}s_{1/2},n_{b}s_{1/2}$ interacting with $n_{\alpha}p_{1/2},n_{\beta}p_{1/2}$ states through the triplet channel with $m=m_{a}+m_{b}=0$. In this case $D=16/9$ and with $n_{a}=99,n_{b}=100,n_{\alpha}=98,n_{\beta}=100$ we find
$V(R)/h=-\frac{44,000}{R^{3}}~{}(\rm MHz)$ (9)
@@
with $R$ in units of $\mu\rm m$. Thus, using $n=100$ Rydberg states at $R=20~{}\mu\rm m$ will give at most about $5~{}\rm MHz$ of interaction strength. This is roughly a factor of five larger than the undressed interaction shown in Fig. 7.
@@
By judicious design of Rydberg pulse profiles it is possible to achieve an effective interaction between the original atoms that is much greater than $V$, although not as large as $V^{\prime}$[165, 166, 167].
@@
We conclude that transversal logical gates with the surface code based on native long-range Rydberg interactions will not scale beyond a code distance of $d=10$ using the surface code, and even that will be extremely challenging to reach. There are methods to increase the Rydberg interaction range based on reduction of the Förster defect, or on adding additional atoms to mediate the interaction. Such methods may extend the interaction range beyond $20~{}\mu\rm m$ which would make a code distance of $d=15$ possible. An alternative path to higher code distance is to use quantum low density parity check (qLDPC) codes that have non-local check operations and a higher code rate $k/n$ than the surface code. Implementation of high rate codes for quantum memory with atom arrays has been analyzed in several recent papers[168, 169, 125, 170]. An optimized neutral atom implementation[125] of the bicycle code introduced in [171] identified a qubit layout for a $d=18$ code that is realizable with long-range Rydberg gates, without invoking atom motion or any improvement from Rydberg dressing.
IV.2 Transport based gates

### v3

@@ 3 Rydberg gate fidelity
The original proposals for entangling pairs of atoms and preparing many-body entanglement using the Rydberg blockade mechanism Jaksch2000 ; Lukin2001 were introduced 25 years ago. The basic ideas and approaches that have been developed for encoding qubits in neutral atoms, their interactions, and mechanisms for performing and characterizing Rydberg quantum gates have been reviewed in many articles Saffman2005a ; Reinhard2007 ; Comparat2010 ; Saffman2010 ; Walker2012 ; JLim2013 ; Marcassa2014 ; Saffman2016 ; Browaeys2016 ; Ryabtsev2017 ; Saffman2019 ; Browaeys2020 ; Henriet2020 ; Adams2020 ; Beterov2020 ; XWu2021 ; Morgado2021 ; Kaufman2021 . First observations of the blockade phenomenon in a many atom ensemble Tong2004 were followed by demonstration of blockade between pairs of individual atoms Gaetan2009 ; Urban2009 . Shortly thereafter the first entanglement and two-qubit quantum gates were achieved using Rydberg interactions Wilk2010 ; Isenhower2010 ; Zhang2010 . The first gate demonstrations had relatively low fidelity and it took some years for the performance to be improved as can be seen in Fig. 4. The improvement can be ascribed to new developments in several directions which together have resulted in recent rapid progress reaching fidelity well over $99\%$ in several experiments. Notably these high fidelity results were achieved with four different atomic elements 87Rb Evered2023 , 133Cs Radnaev2025 , 88Sr Tsai2025 , and 171Yb Peper2025 ; Muniz2025 . While a high fidelity two-qubit gate forms the basis for scaling to larger circuits, in the context of optimizing performance for quantum error correction the two-qubit gate fidelity may not directly translate to the fidelity that is achieved for a syndrome extraction protocol or logical operation. In other words optimizing the gate protocol using the logical performance as a cost function may lead to better performance than optimizing the two-qubit gate in isolation. Several recent papers have addressed this question and predicted improved performance for surface code syndrome extraction Jandura2023 ; Jandura2024b ; FQGuo2025 .
@@
Figure 4: Experimental results for Rydberg $\sf CZ$ gate fidelity. The atomic species are Rb (red circles), Cs (blue squares), Sr (purple triangles), Yb (green diamonds). References: a Wilk2010 , b Isenhower2010 , c Zhang2010 , d Maller2015 , e Jau2016 , f YZeng2017 , g Levine2018 , h Picken2019 , i Levine2019 , j Graham2019 , k Madjarov2020 , l Graham2022 , m ZFu2022 , n Schine2022 , o Evered2023 , p SMa2023 , q King2024 , r Radnaev2025 , s Finkelstein2024 , t ACao2024 , u Muniz2025 , v MChow2024 , w Peper2025 , x Chinnarasu2025 , y Tsai2025 .
@@
One area of progress has been development of improved gate protocols that reduce the amount of time spent in the Rydberg state, thereby decreasing scattering errors, and which are more robust against experimental imperfections Jandura2022 ; Jandura2023 ; Pagano2022 ; Mohan2023 . These time-optimal gates can be designed to work well also without strong blockade Poole2025a , in which case there is some population of doubly-excited Rydberg states during the gate. In this case additional errors arise from inter-atomic forces which must be accounted for Robicheaux2021 .
@@
Another factor contributing to improved performance has been the preparation of colder atoms. Early experiments Wilk2010 ; Isenhower2010 ; Zhang2010 had atom temperatures greater than $50\penalty 10000\ \mu\rm K$ whereas the recent, high fidelity results were all below $10\penalty 10000\ \mu\rm K$. Although the Rydberg blockade interaction is intrinsically insensitive to atom temperature there are higher order effects which lead to errors that grow with temperature. In order to prepare entanglement the atoms must be in a superposition of ground and Rydberg states for some amount of time during the gate sequence. Atomic motion, due either to finite temperature, or photon recoil from the laser excitation Robicheaux2021 , leads to a stochastic gate phase as was first pointed out in Wilk2010 , which degrades the entanglement fidelity. Furthermore as the temperature increases the trapped atoms become progressively more delocalized which leads to fluctuations in the Rydberg interaction strength, as well as excitation errors due to uncertain overlap with focused Rydberg beams. This latter effect can be reduced, but not eliminated, using shaped beams with a flattened intensity profile Reetz-Lamour2008b ; Gillen-Christandl2016 ; Evered2023 ; Chinnarasu2025 .
@@
A technical aspect of the improvement in gate fidelity that deserves highlighting is the role played by laser noise. Even when the atoms are in their motional ground state phase and intensity noise on the excitation lasers contribute to imperfect Rabi oscillations between ground and Rydberg states. Although the issue of laser noise was known in the trapped ion and clock communities Green2013 ; Bishof2013 ; Akerman2015 the specific impact on Rydberg gates with Rabi frequencies in the MHz range had not been studied in detail until the work of Ref. deLeseleuc2018 . There it was shown by numerical simulations that phase noise at the frequency of the ground-Rydberg Rabi frequency was particularly detrimental. Additional analysis established limits on the spectral density of the noise for reaching a desired gate fidelity XJiang2023 ; Day2022 . These limits are not met by the external cavity diode lasers that were typically used for early Rydberg excitation experiments. It was shown convincingly in Levine2018 that cavity filtering of the laser phase noise spectrum led to improved control fidelity and recent demonstration of “spin-locking" diagnostics has enabled direct characaterization of the errors induced by laser noise Tsai2025 .
@@
With all of these advances the state of the art in Rydberg mediated $\sf CZ$ gate fidelity has reached ${\mathcal{F}}=0.9971(5)$ Tsai2025 . It is likely that with further experimental improvements ${\mathcal{F}}>0.999$ will soon be achieved. This may occur first in the alkaline earth atoms Sr and Yb which use a one-photon Rydberg excitation pulse starting from the metastable 3P0 clock state. For the heavy alkali atoms Rb and Cs two-photon excitation from the ground state has been the most widely used approach, although there is also progress on one-photon Rydberg excitation of Cs for implementing an adiabatic dressing gate Keating2015 ; MChow2024 . With two-photon excitation, high fidelity requires extremely low intensity noise in order to stabilize the differential AC Stark shift between ground and Rydberg states Maller2015 that is significant when operating at high intensity and large detuning from the intermediate state used for ladder excitation. High intensity and large detuning are requirements in order to have a fast Rabi rate to minimize Rydberg scattering errors, while also minimizing scattering from the intermediate state.
@@
$P_{\rm R}=\int_{0}^{T}dt\,P_{1}(t)+P_{2}(t)+2P_{12}(t).$ (3)
It was shown in Wesenberg2007 assuming the atomic qubits start in a separable pure state, and considering the growth of the von Neumann entropy during the gate, that in order to create one unit of entanglement we require
@@
It is instructive to estimate this bound using atomic parameters and compare with known gate protocols. As an example consider two Cs atoms excited to the $|68p_{3/2},m_{j}=3/2\rangle$ state that are separated by $R=2.5\penalty 10000\ \mu\rm m$ with the interatomic axis aligned along the quantization axis. The Rydberg intgeraction energy computed with the Pair interaction code Weber2017 is $V=2\pi\times 215\penalty 10000\ \rm MHz$ and the Rydberg lifetime in a room temperature environment is Beterov2009 ; Beterov2009b $\tau_{\rm R}=198\penalty 10000\ \mu\rm s$. These values give
@@
There is an important open question of how to design $\Omega(t)$ to reach this bound. The original Rydberg $\pi-2\pi-\pi$ gate proposal Jaksch2000 has an error scaling XZhang2012 $\epsilon\sim 1/(V\tau_{\rm R})^{2/3}$ which is found from optimizing the Rabi frequency to balance errors between imperfect rotations at finite interaction strength ( $\sim|\Omega/V|^{2}$) and the Rydberg scattering error $(\sim 1/\Omega\tau_{\rm R}).$ It turns out that the $\pi-2\pi-\pi$ gate can be modified to have an error scaling as $1/(V\tau_{\rm R})$ Cole2023 , and also other protocols that rely on individual addressing of control and target qubits reach the optimum $1/(V\tau_{\rm R})$ scaling Petrosyan2017 .
@@
Nevertheless, protocols that have $\Omega_{1}(t)\neq\Omega_{2}(t)$ so the temporal waveforms are different on the two qubits require more complex control systems. Adiabatic proposals for gates with symmetric qubit control $(\Omega_{1}(t)=\Omega_{2}(t))$ have been proposed YSun2020 ; Saffman2020 , and the qubit symmetric, two-pulse protocol from Levine2019 was the first to demonstrate fidelity above 97%. The time-optimal protocols Jandura2022 ; Pagano2022 ; Mohan2023 which have delivered the highest gate fidelities are qubit symmetric but do not rely on adiabaticity to return all initial states to the ground state at the end of the gate. Instead they rely on fast driving to minimize time spent in the Rydberg state with a time dependent phase profile that ensures that initial states with both one- and two-atoms Rydberg coupled completely return to the ground state.
@@
In the limit of strong blockade the protocols that come closest to the limit set by Eq. (5) are the two-atom dark state gate Petrosyan2017 which reaches $\epsilon/\epsilon_{\rm min}\simeq 19$, the time-optimal gate Jandura2022 which reaches $\epsilon/\epsilon_{\rm min}\simeq 15$ and the modified $\pi-2\pi-\pi$ gate Cole2023 . With strong blockade additional errors due to simultaneous Rydberg excitation of both atoms, which leads to inter-atomic forces, are negligible. In the limit of weak blockade modified time-optimal profiles were found to reach $\epsilon/\epsilon_{\rm min}=2.1$ Poole2025a . Including two-atom interaction dephasing and the effect of photon recoils the predicted error increases to $\epsilon/\epsilon_{\rm min}=3.0$. Photon recoil heating can be minimized using more complex time-dependent pulses Lis2023 ; ZZhang2024 ; vanDamme2025 , but this has not yet been demonstrated in practice with Rydberg gates.
4 The connectivity challenge
@@
Running complex quantum algorithms on a digital quantum computer requires gate operations involving qubits that may not be next to each other. Any physical interaction between qubits has a finite range. While there are settings for which all qubits are interconnected, such cases do not support scaling to arbitrarily large qubit numbers. As an example consider coupling many atomic qubits to the mode of an optical resonator such that photons mediate all to all connectivity Pellizzari1995 ; Vaidya2018 ; Ramette2022 . Even though photon mediated coupling does not intrinsically have a limited range there is a practical limit to how many atoms can be strongly coupled to a single cavity.
@@
For quantum computation without error correction the additional gate operations needed to connect non-proximal qubits add operational overhead and reduce the useful computational depth. The severity of establishing connectivity depends on the algorithm being run, and can be reduced with optimized compilation that is aware of the physical level hardware capabilities DBTan2024 ; Schmid2024 . There are also encoding techniques for specific problems that effectively provide long-range interactions at the cost of consuming more physical qubits Lechner2015 ; Nguyen2023 . A well designed error corrected processor that operates with logical qubits should have a computational depth that greatly exceeds that of a NISQ machine. Nevertheless, useful error correction does not imply an arbitrarily small logical error rate, just an error rate that is smaller than that of the underlying physical operations. Thus also logical processors will suffer reduced useful circuit depth due to resources required for connectivity, as well as longer operation times.
In this section we compare several approaches to establishing non-local connectivity for logical qubits in atom arrays with an emphasis on the execution time requirements. Performing a gate between neighboring logical qubits may require $\it nonlocal$ connectivity at the physical level when the gate is based on coherent interactions. An important alternative is lattice surgery which enables gates between logical qubits, requiring only local physical gates plus measurements and auxiliary qubit resources Horsman2012 . To focus the discussion we will assume encoding in the $d^{2},1,d]]$ rotated surface code. This encoding requires $2d^{2}-1$ physical qubits for each logical qubit assuming no reuse of ancillas between $\sf X$ and $\sf Z$ stabilizer extractions. Based on a standard depolarizing noise model the logical error rate of the rotated code with a minimum-weight perfect-matching decoder is ORourke2025
$p_{\rm L}=0.08(p/p_{\rm th})^{0.58d-0.27}.$ (7)
where the threshold physical error rate is $p_{\rm th}=0.0053.$ Taking ${\mathcal{F}}=0.9992$ or $p=0.0008$, which appears within reach of Rydberg gates, a code distance of $d=10$ is sufficient to reach within about a factor of two of $1/p_{\rm L}=10^{6}$. At this code distance 100 logical qubits would require 19900 physical qubits, plus the overhead of magic state distillation for non Clifford operations. Such a “Megaquop" machine with 100 error corrected logical qubits is likely to be of great interest from a scientific perspective Preskill2025 . While atom arrays are currently one of the most promising approaches for reaching this combination of low error rates and high qubit counts the clock rate of neutral atoms is slow compared to other platforms. With that challenge in mind the following sections provide a comparison of logical gate times based on several different approaches.
4.1 long-range Rydberg interactions
The rotated surface code has $n,k,d]]=[[d^{2},1,d]]$ and the gate time for a transversal logical $\sf CZ$ using sequential long-range Rydberg interactions is $t_{\rm gate}\leq d^{2}(t_{\sf CZ}+t_{\rm beam})$ where $t_{\sf CZ}$ is the time for the Rydberg gate and $t_{\rm beam}$ is the time to point the optical control beams at a new atom pair. This is a strict equality for sequential gate operations and can be quicker if some of the atom pairs are far enough separated to allow for parallel gate operations (see Poole2025a for an analysis of the separation required to minimize interference effects between gate pairs). Significantly, the ability to implement $\sf CZ$ gates transversally, in contrast to the lattice surgery methods discussed in Sec. 4.3, allows for efficient decoding with a reduced number of measurement rounds HZhou2025 .
Figure 6 shows the logical gate time assuming $d=10$. The calculated times which can be well under $100\penalty 10000\ \mu\rm s$ assume two-qubit Rydberg gates with an optical control system that excites pairs of atoms using dynamically scanned, tightly focused beams Graham2022 ; Radnaev2025 . The achievable gate speed is strongly dependent on the performance of the optical beam scanning device. Acousto-optic deflectors are compatible with the sub-microsecond times assumed in the figure, whereas other technologies such as spatial light modulators and digital mirror arrays operate at much slower speeds. The combination of these approaches may allow fast reconfiguration together with the ability to address large qubit arrays Graham2023 .
A limitation of this approach is the finite range of the Rydberg interaction. At $d=10$ and a tight spacing between physical qubits of $r=1.4\penalty 10000\ \mu\rm m$ the required interaction distance is $R=dr=14\penalty 10000\ \mu\rm m$ which is beyond the range of demonstrated Rydberg gates. A high fidelity gate at this distance appears feasible by combining progress in several directions. The use of highly excited Rydberg states can provide close to $V=2\pi\times 6\penalty 10000\ \rm MHz$ of interaction strength at this distance as shown in Fig. 7. With this value for $V$ a modified time-optimal gate profile that provides high fidelity with modest interaction strength can reach ${\mathcal{F}}=0.9992$ with a gate time of $0.46\penalty 10000\ \mu\rm s$ Poole2025a . With an optical beam scanning time of $0.5\penalty 10000\ \mu\rm s$ this leads to a logical gate time of about $90\penalty 10000\ \mu\rm s$.
Extending this approach to larger code distance which implies larger $R$ appears challenging. Our assumption of $r=1.4\penalty 10000\ \mu\rm m$ atomic spacing cannot be significantly reduced without running in to the limit set by the wavefunction of a Rydberg excited electron overlapping with a neighboring ground state atom. The Rydberg interaction strength scales as $1/R^{3}$ at short distances (limit of resonant dipolar interactions) and $1/R^{6}$ at long-range(van der Waals limit) Walker2008 . Judicious application of dc or microwave fields can be used to reduce the Förster defects and extend the $1/R^{3}$ scaling to longer distances Bohlouli-Zanjani2007 ; Ravets2014 ; Beterov2016b ; Kurdak2025 .
Figure 7: Rydberg interaction strength for Cs atoms excited to $100s_{1/2},m_{j}=1/2$ states quantized along the atomic separation axis ($\hat{z}$) with a magnetic field of $B_{z}=10^{-5}\penalty 10000\ \rm T$ . Calculation performed with pair interaction code Weber2017 .
It is perhaps interesting to check how large the interaction could ideally be at say $R=20\penalty 10000\ \mu\rm m$ which could enable $d=15$ and 100 Megaquop performance. At Förster resonance the interaction of two atoms excited to Rydberg states that interact through a single channel can be expressed as $V(R)=-C_{3}D^{1/2}/R^{3}$ where $D$ is an angular factor depending on the interaction channel and
@@
where $e$ is the electronic charge, $\epsilon_{0}$ is the permittivity of free space, the reduced matrix elements are calculated in the fine structure basis, and $\gamma$ denotes the quantum numbers of the initial Rydberg levels $a,b$ that are coupled to levels $\alpha,\beta$ through a Förster interaction Walker2008 ; Beterov2015 . Consider as an example Cs $a,b$ states $n_{a}s_{1/2},n_{b}s_{1/2}$ interacting with $n_{\alpha}p_{1/2},n_{\beta}p_{1/2}$ states through the triplet channel with $m=m_{a}+m_{b}=0$. In this case $D=16/9$ and with $n_{a}=99,n_{b}=100,n_{\alpha}=98,n_{\beta}=100$ we find
$V(R)/h=-\frac{44,000}{R^{3}}\penalty 10000\ (\rm MHz)$ (9)
@@
with $R$ in units of $\mu\rm m$. Thus, using $n=100$ Rydberg states at $R=20\penalty 10000\ \mu\rm m$ will give at most about $5\penalty 10000\ \rm MHz$ of interaction strength. This is roughly a factor of five larger than the undressed interaction shown in Fig. 7.
@@
By judicious design of Rydberg pulse profiles it is possible to achieve an effective interaction between the original atoms that is much greater than $V$, although not as large as $V^{\prime}$ Cesa2017 ; YSun2024 ; Doultsinos2025 .
@@
We conclude that transversal logical gates with the surface code based on native long-range Rydberg interactions will not scale beyond a code distance of $d=10$ using the surface code, and even that will be extremely challenging to reach. There are methods to increase the Rydberg interaction range based on reduction of the Förster defect, or on adding additional atoms to mediate the interaction. Such methods may extend the interaction range beyond $20\penalty 10000\ \mu\rm m$ which would make a code distance of $d=15$ possible. An alternative path to higher code distance is to use quantum low density parity check (qLDPC) codes that have non-local check operations and a higher code rate $k/n$ than the surface code. Implementation of high rate codes for quantum memory with atom arrays has been analyzed in several recent papers Viszlai2023 ; QXu2024 ; Poole2025a ; Pecorari2025 . An optimized neutral atom implementation Poole2025a of the bicycle code introduced in Bravyi2024 identified a qubit layout for a $d=18$ code that is realizable with long-range Rydberg gates, without invoking atom motion or any improvement from Rydberg dressing.
4.2 Transport based gates